Consider the following information for a period of years: Arithmetic Mean Long-term government bonds 6.2 % Long-term corporate bonds 6.3 Inflation 3.3 a. What is the real return on long-term government bonds? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What is the real return on long-term corporate bonds? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a.

Real Return on Government bonds = 0.062 - 0.033 = 2.90%

b.

Real Return on Long-term corporate bonds = 0.063 - 0.033 = 3.00%